OpenStudy (anonymous):

When you simplify each side of the equation 2 - 7(3n + 2) = -5(4n + 2), you will get the equivalent equation -15n - 10 = -20n - 10. TRUE FALSE

OpenStudy (anonymous):

I believe its true

OpenStudy (anonymous):

i believe its false. clear out the brackets first, before adding or subtracting

OpenStudy (anonymous):

\[2-7(3n+2)=-5(4n+2)\]\[2-21n-14=-20n-10\]\[-21n-12=-20n-10\] False

OpenStudy (anonymous):

Thanks guys